OpenStudy (anonymous):

40 is what percent of 64?

OpenStudy (anonymous):

40/64*100 =62.5%

OpenStudy (turingtest):

40=(x/100)64

OpenStudy (turingtest):

man that took forever to delete...

OpenStudy (anonymous):

x=40/6400 %